# ParGo: Bridging Vision-Language with Partial and Global Views

### v1

@@ Abstract
This work presents ParGo, a novel Partial-Global projector designed to connect the vision and language modalities for Multimodal Large Language Models (MLLMs). Unlike previous works that rely on global attention-based projectors, our ParGo bridges the representation gap between the separately pre-trained vision encoders and the LLMs by integrating global and partial views, which alleviates the overemphasis on prominent regions. To facilitate the effective training of ParGo, we collect a large-scale detail-captioned image-text dataset named ParGoCap-1M-PT, consisting of 1 million images paired with high-quality captions. Extensive experiments on several MLLM benchmarks demonstrate the effectiveness of our ParGo, highlighting its superiority in aligning vision and language modalities. Compared to conventional Q-Former projector, our ParGo achieves an improvement of 259.96 in MME benchmark. Furthermore, our experiments reveal that ParGo significantly outperforms other projectors, particularly in tasks that emphasize detail perception ability. The code and models are released at https://github.com/AlanWang0o0/ParGo
@@ Introduction
footnotetext: * Equal contribution. $\dagger$ This work is done when An-Lan Wang is an intern at ByteDance China.
Recent Multi-Modal Large Language Models (MLLMs) (OpenAI 2023b; Team et al. 2023; Liu et al. 2023b; Li et al. 2022) achieve remarkable progress across various tasks (e.g., Visual Question Answering (VQA)). The vision-language projector as a widely used component in MLLMs, aims to provide LLMs with proper visual features. Due to its critical role in bridging modalities, it has garnered significant attention in recent research (Cha et al. 2023; Alayrac et al. 2022; Zhu et al. 2023).
@@
Figure 1: Illustration of the global and partial information. An image can be properly described by the two kinds of information. Globally, this image shows a landscape of an ancient stone fortress. Delve into the partial information, two individuals stand atop the fortress, the wooden gate at the bottom of the fortress is partially open, and so forth. Please zoom in and view in color.
@@
Large Language Models (LLMs) (Touvron et al. 2023; OpenAI 2023b; Chiang et al. 2023) have achieved remarkable progress, leading recent works (Alayrac et al. 2022; Team et al. 2023; Bai et al. 2023) to generalize this success to more modalities, i.e.,, Multimodal Large Language Models (MLLMs). In those works, closed-source works (OpenAI 2023b; Team et al. 2023; Bai et al. 2023) have shown great advancements, highlighting their high performance on complex tasks. In contrast, open-source models have also made significant progress, promoting transparency and collaboration in the research community. Pioneer works (Li et al. 2023b; Alayrac et al. 2022) established competitive baselines by integrating massive image-text pairs. Furthermore, recent works (Liu et al. 2023b; Dai et al. 2023) (Liu et al. 2023a, b; Chen et al. 2023b) boosting the zero-shot capabilities on various downstream tasks via collecting more high-quality multi-modal instruction (visual instruction tuning). On the other hand, recent works (Liu et al. 2024a; Hu et al. 2023; Liu et al. 2024b) also focus on fine-grained understanding (e.g., text recognition).
Figure 2: (a). The pipeline of a MLLM with our proposed ParGo as the vision-language projector. First of all, we use a frozen image encoder to extract image features. To better align the pre-trained visual encoder with the LLM, we propose a Partial-Global projector to project the image features using two kinds of tokens i.e., partial and global tokens. Finally, the output partial and global visual tokens, as well as the tokenized text, are fed into the LLM to generate the text output in an auto-regressive manner. Specifically, each Partial-Global projector layer contains a Partial-Global Perception block that utilizes two kinds of tokens, i.e., partial and global tokens, to extract the image features. Additionally, to fully consider the relation between different partial regions in an image, a cascaded partial perception block is incorporated to enable interactions between partial tokens in a cascaded manner. (b). A Demonstration of the Partial-Global and the Cascaded Partial Attention mask. It’s worth noting that the Partial-Global Attention mask remains the same in different layers, while the Cascaded Partial Attention mask changes across various layers.
@@
Training models using web-crawled large-scale image-text datasets (e.g., (Schuhmann et al. 2021; Byeon et al. 2022; Changpinyo et al. 2021; Sharma et al. 2018)) has become the most common strategy for MLLMs. Nevertheless, web-crawled datasets primarily present the main feature of the image using noisy and short captions, lacking detailed descriptions. To obtain detailed descriptions, some works provide boxes (or mask) level captions but are constrained by the box-generation model. The recent remarkable progress achieved by close-sourced MLLMs (OpenAI 2023b; Team et al. 2023) has led recent researchers(Chen et al. 2023b; Yu et al. 2024a) to consider using MLLM to synthesize detail-captioned data, supplementing the limitations of conventional web-crawled datasets. In this work, we further contribute a detail-captioned dataset for pre-training, aimed at enhancing the alignment between the two modalities from a data perspective.
@@
For the supervised fine-tuning stage, following previous work (Cha et al. 2023), we employ four types of tasks, including: 1) Open-ended VQA, i.e. VQAv2 (Goyal et al. 2017), GQA (Hudson and Manning 2019), OCRVQA (Mishra et al. 2019), VSR (Liu, Emerson, and Collier 2023). 2) Multiple-Choice VQA, including ScienceQA (Lu et al. 2022), A-OKVQA (Schwenk et al. 2022). 3) Referring Expression Comprehension (REC), which includes RefCOCO (Kazemzadeh et al. 2014) and VG (Krishna et al. 2017). 4) Instruction tuning data, LLaVA150k (Liu et al. 2023b). For each dataset, we use the same templates as previous works (Liu et al. 2023a; Cha et al. 2023).
@@
In Figure 3, we give several examples from the MM-Vet benchmark to demonstrate the superiority of our proposed Partial-Global projector in preserving partial and global information. Specifically, we compare our full model with the baseline that uses a Q-Former with 304 tokens as the projector. From the first two examples (in the first column), we find that our model has a better perception of the location-specified partial object (i.e., the person in the front left, the green logo on the car.), and correctly answers the question. Conversely, due to the Q-Former projector’s overemphasis on prominent regions and neglect of image details, the baseline fails to answer the question correctly, e.g., the blue Ford logo in the car. Regarding the two optical character recognition examples in the middle column, our model accurately identifies the characters in the image and correctly understands the text. In contrast, the baseline model fails to achieve correct recognition in these instances. The two examples in the last column illustrate that our method has a better perception of the overall images while maintaining the partial information, such as the colors of different cats and the presence of the elephant on the shoreline. In summary, the qualitative results highlight the ability of our proposed Projector to align the two modalities, providing the LLM with features covering both partial and global information.
@@ Conclusion
In this work, we focus on the vision-language projector in MLLMs, proposing Partial-Global projector (ParGo). ParGo employs partial and global tokens with specially designed attention masks to extract two kinds of information separately, with considering the relation between different partial regions in an image. Moreover, to further facilitate the alignment between the two modalities, we contribute a large-scale detail-captioned dataset ParGoCap-1M-PT for pre-training. Extensive ablations and experiments are conducted, which illustrate the effectiveness of our ParGo. For instance, our ParGo outperforms Q-Former by 259.96 scores in the MME benchmark. We find that ParGo significantly outperforms other projectors, particularly in tasks that emphasize detail perception. These results highlight ParGo’s potential to enhance MLLMs by providing a more nuanced understanding of visual content through the integration of both partial and global views.

### v2

@@ Abstract
This work presents ParGo, a novel Partial-Global projector designed to connect the vision and language modalities for Multimodal Large Language Models (MLLMs). Unlike previous works that rely on global attention-based projectors, our ParGo bridges the representation gap between the separately pre-trained vision encoders and the LLMs by integrating global and partial views, which alleviates the overemphasis on prominent regions. To facilitate the effective training of ParGo, we collect a large-scale detail-captioned image-text dataset named ParGoCap-1M-PT, consisting of 1 million images paired with high-quality captions. Extensive experiments on several MLLM benchmarks demonstrate the effectiveness of our ParGo, highlighting its superiority in aligning vision and language modalities. Compared to conventional Q-Former projector, our ParGo achieves an improvement of 259.96 in MME benchmark. Furthermore, our experiments reveal that ParGo significantly outperforms other projectors, particularly in tasks that emphasize detail perception ability.
Code and modelshttps://github.com/bytedance/ParGo
@@ Introduction
Recent Multi-Modal Large Language Models (MLLMs) (OpenAI 2023b; Team et al. 2023; Liu et al. 2023b; Li et al. 2022) achieve remarkable progress across various tasks (e.g., Visual Question Answering). The vision-language projector as a widely used component in MLLMs, aims to provide LLMs with proper visual features. Due to its critical role in bridging modalities, it has garnered significant attention in recent research (Cha et al. 2023; Alayrac et al. 2022; Zhu et al. 2023).
@@
Figure 1: Illustration of the global and partial information. An image can be properly described by the two kinds of information. Globally, this image shows a landscape of an ancient stone fortress. Delve into the partial information, two individuals stand atop the fortress, the wooden gate at the bottom of the fortress is partially open, and so forth.
@@
Large Language Models (LLMs) (Touvron et al. 2023; OpenAI 2023b; Chiang et al. 2023) have achieved remarkable progress, leading recent works (Alayrac et al. 2022; Team et al. 2023; Bai et al. 2023) to generalize this success to more modalities, i.e.,, Multimodal Large Language Models (MLLMs). In those works, closed-source works (OpenAI 2023b; Team et al. 2023; Bai et al. 2023) have shown great advancements, highlighting their high performance on complex tasks. In contrast, open-source models have also made significant progress, promoting transparency and collaboration in the research community. Pioneer works (Li et al. 2023b; Alayrac et al. 2022; Shan et al. 2022a, b) established competitive baselines by integrating massive image-text pairs. Furthermore, recent works (Liu et al. 2023b; Dai et al. 2023) (Liu et al. 2023a, b; Chen et al. 2023b) boosting the zero-shot capabilities on various downstream tasks via collecting more high-quality multi-modal instruction. On the other hand, recent works (Liu et al. 2024a; Hu et al. 2023; Liu et al. 2024b; Zhao et al. 2024a, b; Liu et al. 2023d; Li et al. 2024; Lin et al. 2024) also focus on fine-grained understanding (e.g., text recognition, action understanding).
Figure 2: (a). The pipeline of a MLLM with our proposed ParGo as the vision-language projector. First of all, we use a frozen image encoder to extract image features. To better align the pre-trained visual encoder with the LLM, we propose a Partial-Global projector to project the image features using two kinds of tokens i.e., partial and global tokens. Finally, the output partial and global visual tokens, as well as the tokenized text, are fed into the LLM to generate the text output in an auto-regressive manner. Specifically, each Partial-Global projector layer contains a Partial-Global Perception block that utilizes two kinds of tokens to extract the image features. Additionally, to fully consider the relation between different partial regions in an image, a cascaded partial perception block is incorporated to enable interactions between partial tokens in a cascaded manner. (b). A Demonstration of the Partial-Global and the Cascaded Partial Attention mask. It’s worth noting that the Partial-Global Attention mask remains the same in different layers, while the Cascaded Partial Attention mask changes across various layers.
@@
Training models using web-crawled large-scale image-text datasets (e.g., (Schuhmann et al. 2021; Byeon et al. 2022; Changpinyo et al. 2021; Sharma et al. 2018)) has become the most common strategy for MLLMs. Nevertheless, web-crawled datasets primarily present the main feature of the image using noisy and short captions, lacking detailed descriptions. To obtain detailed descriptions, some works (Wang et al. 2023, 2025) provide boxes (or mask) level captions but are constrained by the box-generation (grounding) model (Fang et al. 2024, 2025; Tang et al. 2022a, b). The recent remarkable progress achieved by close-sourced MLLMs has led recent researchers (Chen et al. 2023b; Yu et al. 2024a) to consider using MLLM to synthesize detail-captioned data, supplementing the limitations of conventional web-crawled datasets. In this work, we further contribute a detail-captioned dataset for pre-training, aimed at enhancing the alignment between the two modalities from a data perspective.
@@
For the supervised fine-tuning stage, following previous work (Cha et al. 2023), we employ four types of tasks, including: 1) Open-ended VQA, i.e. VQAv2 (Goyal et al. 2017), GQA (Hudson and Manning 2019), OCRVQA (Mishra et al. 2019), VSR (Liu, Emerson, and Collier 2023). 2) Multiple-Choice VQA, including ScienceQA (Lu et al. 2022), A-OKVQA (Schwenk et al. 2022). 3) Referring Expression Comprehension (REC), which includes RefCOCO (Kazemzadeh et al. 2014) and VG (Krishna et al. 2017). 4) Instruction tuning data, LLaVA150k (Liu et al. 2023b). For each dataset, we use the same templates as previous works (Liu et al. 2023a; Cha et al. 2023; Shan et al. 2024).
@@
In Figure 3, we give several examples to demonstrate the superiority of our proposed Partial-Global projector in preserving partial and global information. Specifically, we compare our full model with the baseline that uses a Q-Former with 304 tokens as the projector. From the first two examples (in the first column), we find that our model has a better perception of the location-specified image part (i.e., the person in the front left, the green logo on the car.). Conversely, due to the Q-Former projector’s overemphasis on prominent regions and neglect of image details, it fails to answer the question correctly. Regarding the two examples in the middle column, our model accurately identifies the characters in the image and correctly understands the text. In contrast, the Q-Former fails to achieve correct recognition in these instances. The last two examples illustrate that our method has a better perception of the overall images while maintaining the partial information, i.e., the colors of different cats and the presence of the elephant on the shoreline. In summary, the qualitative results highlight the ability of our proposed Projector to align the two modalities, providing the LLM with features covering both partial and global information.
@@ Conclusion
In this work, we focus on the vision-language projector in MLLMs, proposing Partial-Global projector (ParGo). ParGo employs partial and global tokens with specially designed attention masks to extract two kinds of information separately, with considering the relation between different partial regions in an image. Moreover, to further facilitate the alignment between the two modalities, we contribute a large-scale detail-captioned dataset ParGoCap-1M-PT for pre-training. Extensive ablations and experiments are conducted, which illustrate the effectiveness of our ParGo. We find that ParGo significantly outperforms other projectors, particularly in tasks that emphasize detail perception. These results highlight ParGo’s potential to enhance MLLMs by providing a more nuanced understanding of visual content through the integration of both partial and global views.
Acknowledgments
This work was partially supported by the National Key Research and Development Program of China (2023YFA1008503), NSFC(92470202, U21A20471), and Guangdong NSF Project (No. 2023B1515040025).